### v1

# I Know About “Up”! Enhancing Spatial Reasoning in Visual Language Models Through 3D Reconstruction

Abstract
Visual Language Models (VLMs) are essential for various tasks, particularly the visual reasoning tasks, due to their robust multi-modal information integration, visual reasoning capabilities, and contextual awareness. However, existing VLMs’ visual spatial reasoning capabilities are often inadequate, struggling even with basic tasks such as distinguishing left from right. To address this, we propose the ZeroVLM [^1] Code: https://github.com/zhouhao028/Iknow_up model, designed to enhance the visual spatial reasoning abilities of VLMs. ZeroVLM employs Zero-1-to-3, a 3D reconstruction model for obtaining different views of the input images and incorporates a view prompt to further improve visual spatial reasoning. Experimental results on four visual spatial reasoning datasets show that our ZeroVLM achieves up to 19.48% accuracy improvement, which indicates the effectiveness of 3D reconstruction and view prompt of our ZeroVLM.
I Know About “Up”! Enhancing Spatial Reasoning in Visual Language Models Through 3D Reconstruction
Hao Zhou Guangdong Polytechnic Normal University hao.zhou28@outlook.com                        Zaiqiao Mengthanks: Corresponding Author. University of Glasgow zaiqiao.meng@glasgow.ac.uk                        Yifang Chen Guangdong Polytechnic Normal University chenyf@gpnu.edu.cn
1 Introduction
Visual Language Models (VLMs), such as LLaVA (Liu et al., 2024), MiniGPT-4 (Zhu et al., 2023) and InternGPT (Liu et al., 2023d), are a class of deep neural models adept at simultaneously processing and understanding both visual and linguistic information. These VLMs often consist of an image encoder, an embedding projector to align image and text representations and a text decoder to process the projected image embedding and text representations, enabling joint understanding and reasoning between these modalities (Li et al., 2022). By using the language generation power of underlining text decoders, these VLMs showcase remarkable interaction capabilities in various applications, such as referring expression comprehension (Li et al., 2019b; Yu et al., 2016), visual question answering (Goyal et al., 2017; Zellers et al., 2019; Hudson and Manning, 2019), visual language reasoning (Liu et al., 2022a), and entailment (Liu et al., 2021; Xie et al., 2019).
Figure 1: An example of the VQA task, where humans can easily recognize positions under different views, but the vanilla LLaVA (Liu et al., 2024) can only predict correctly from certain views. By performing 3D reconstruction to obtain different views of the image, we can improve LLaVA’s predictive accuracy.
However, many visual language tasks, such as visual question answering (VQA) (Redmon et al., 2016; Long et al., 2015; Zhou et al., 2017; Dai et al., 2018; Hu et al., 2019) and image segmentation (Huang et al., 2023), require the ability to recognize spatial information from images. For instance, in visual question answering, a system must understand the spatial relationships between objects to correctly answer questions about an image (Dai et al., 2018; Hu et al., 2019). Visual spatial reasoning in VLMs requires multi-modal understanding (Khattak et al., 2023), cross-modal mapping (Wang et al., 2023), inference of visual spatial relations (Liu et al., 2023a), and integration of context comprehension (Zhao et al., 2023). This is because VLMs must comprehend intricate spatial relationships within images, which are hierarchical and multifaceted, involving various relationships among multiple objects (such as containment, proximity, overlap, etc.) (Chen et al., 2024). Enhancing visual spatial reasoning capabilities of models can significantly elevate their performance in image comprehension and processing (Yang et al., 2024).
Presently, numerous studies explore various benchmarks and methods to enhance the visual spatial reasoning capabilities of VLMs to improve their performance, exemplified by visual spatial reasoning (Liu et al., 2023a) and Whatsup_vlm (Kamath et al., 2023). These endeavors employ diverse methodologies (Su et al., 2019; Qi et al., 2020; Kiela et al., 2019) to augment the visual spatial reasoning abilities of models. However, despite these efforts yielding improvements in model performance to some extent, substantial challenges persist in comprehending complex scenes and intricate spatial relationships. Existing methodologies predominantly rely on 2D information, impeding the comprehensive capture of 3D spatial relationships among objects. For example, Figure 1 illustrates an example of the VQA task. While humans could easily recognize positions under different views of the objects, LLaVA can only predict correctly from certain views due to the lack of image views during their pretraining.
To tackle this challenge, we propose a novel model, called ZeroVLM, to solve the visual spatial reasoning task through the 3D reconstruction technique, entailing the reconstruction of images in 3D and capturing them from various views. In particular, our ZeroVLM utilize the Zero-1-to-3 (Liu et al., 2023c) model for constructing 3D views from a single 2D image within our tested datasets. This allows VLMs to access richer spatial information through 3D transformation, thereby enhancing their visual spatial reasoning capabilities. We conducted comprehensive experiments across four visual spatial reasoning datasets, and the results show that all the tested VLMs have notably improved spatial reasoning capabilities through our ZeroVLM.
Our contributions are summarized as follows:
  • We proposed the ZeroVLM model, a new model for visual spatial reasoning tasks, which can generate images of different views based on the original image to enhance spatial reasoning ability.
  • Our ZeroVLM utilises a 3D reconstruction model to obtain different views from the input image, which enhances spatial relationships at the image level to improve the visual spatial reasoning capabilities of VLMs.
  • We validate the effectiveness of our ZeroVLM through visual spatial reasoning experiments conducted on four different datasets.
(a) Single-View
(b) Multi-View
Figure 2: Single-view images were generated using Zero-1-to-3 to produce left-view, right-view, and random-view images. Multi-view images were created by combining these different single-view images in various configurations.
2 Related Work
Vision Language Models (VLMs). VLMs amalgamate computer vision and natural language processing technologies to comprehend and generate correlations between images and natural language (Dosovitskiy et al., 2020). VLMs can accept image input to generate corresponding response in text (Menon and Vondrick, 2022). In terms of training architecture, recent works include joint pretraining architectures (e.g. OFA Wang et al. (2022) and VLMo (Bao et al., 2021)), which train image and text data jointly, and image-to-text mapping architectures (e.g., PaLI (Chen et al., 2022b)), which map an image encoder to a well-pretrained text encoder. The latter approach has gained more popularity. VLMs exhibit cross-modal understanding capabilities, enabling them to extract information from images and translate it into natural language, or retrieve information from natural language and generate corresponding images (Chen et al., 2022a). This wide range of capabilities broadens the potential of VLMs (e.g., GlaMM (Rasheed et al., 2023) and MiniGPT-4 (Zhu et al., 2023)) in understanding and generating associations between images and language, such as Image Captioning (Zhang et al., 2021), Visual Question Answering (Yang et al., 2022), and Multi-modal Translation (Li et al., 2019a). In this study, we aim to explore the visual-spatial recognition capabilities of these VLMs.
Visual Spatial Reasoning. Visual spatial reasoning refers to the cognitive ability to understand and manipulate the spatial relationships of objects or elements in a given environment (Liu et al., 2023a). This reasoning ability involves not only recognizing the properties of individual entities, but also understanding the complex relationships and structures between them. Recent works have aimed to benchmark the problem of visual spatial reasoning, notably VSR (Liu et al., 2023a) and Whatsup_vlm (Kamath et al., 2023). VSR (Liu et al., 2023a) evaluates the ability of VLMs using text-image pairs to describe various visual spatial relationships, while Whatsup_vlm (Kamath et al., 2023) assesses the visual spatial reasoning of VLMs through specific prepositions and perspectives. Although these methods provide valuable insights into the visual spatial reasoning capabilities of VLMs, they remain somewhat limited. In our work, we employ 3D reconstruction to more comprehensively test the visual spatial reasoning ability of VLMs.
3D Reconstruction. Visual spatial reasoning encompasses the cognitive ability to comprehend and manipulate visual spatial information, while 3D reconstruction involves generating a three-dimensional representation of an object or scene from two-dimensional images or other sensor data. For instance, MonoScene (Cao and De Charette, 2022) infers dense geometric structures and semantic information of a scene from a single monocular RGB image, offering an efficient and innovative approach to complete 3D semantic scenes. VoxFormer (Li et al., 2023) proposes a transformer-based (Vaswani et al., 2017) semantic scene completion framework, addressing camera-based 3D semantic scene completion by introducing sparse voxel queries (Guchait et al., 2021) and masked autoencoder design (He et al., 2022). Zero-1-to-3 (Liu et al., 2023c) is a 3D reconstruction model that excels in zero-shot generalization on out-of-distribution datasets and unprocessed images by using a single image as input without requiring additional 3D or depth information. We chose this model for its simplicity and effectiveness compared to MonoScene (Cao and De Charette, 2022) and VoxFormer (Li et al., 2023).
Figure 3: An overview of our proposed ZeroVLM model. Our ZeroVLM first uses Zero-1-to-3 to perform 3D reconstruction to obtain different views of the input image, and then it stitches the original images with these different views to obtain the stitched image, which is the input of a VLM for answer prediction.
3 Preliminaries
In this section, we will formally define the task of visual spatial reasoning and introduce the Zero-1-to-3 model (Liu et al., 2023c), which is one of the components utilized by our model.
3.1 Visual Spatial Reasoning Task
Visual spatial reasoning is critically important in the fields of computer vision, artificial intelligence, and machine learning (Zakari et al., 2022; Rajabi and Kosecka, 2023; Liu et al., 2022b). The visual spatial reasoning task involves inferring and understanding objects, scenes, or spatial relationships, addressing location, orientation, and spatial arrangement for accurate object localization and recognition. This task is normally formulated as a visual question answering task (Liu et al., 2021; Kamath et al., 2023) by asking and answering questions about objects on images and through context in text. Formally, we denote an image and its associated question as i and $q$, respectively. The questions in this task typically ask to identify the relationships among multiple objects or to locate specified objects within the image. Figure 1 shows an example of this task. The goal of this task is to answer this question by accurately comprehending the visual spatial relations and discerning the visual spatial position of target objects relative to others.
3.2 Zero-1-to-3
Zero-1-to-3 (Liu et al., 2023c) is a model designed for visual reconstruction tasks. It works by identifying objects in input image i and adjusting the camera perspective of these objects. Given a image $i\in\mathbb{R}^{H\times W\times 3}$ and a relative camera rotation matrix $R\in\mathbb{R}^{3\times 3}$ and a translation vector $T\in\mathbb{R}^{3}$, Zero-1-to-3 leans a model $f$ such that its output is a new perspective image under the specified camera transformation:
$\hat{i}_{R,T}=f(i,R,T)$ (1)
where $\hat{i}_{R,T}$ represents the generated new perspective image. When we input original image i and hope to generate a 3D reconstructed view from the left at an angle of 45° through Zero-1-to-3, we can achieve this effect by setting $R$ to the rotation matrix of 45° around the Y axis and setting $T$ to the vector translated to the left by a certain distance. In order to synthesize new views under such partial constraints, Zero-1-to-3 uses a large-scale diffusion model to integrate the geometric prior knowledge of the input image. The conditional diffusion model embedded in Zero-1-to-3  (Yang et al., 2023; Podell et al., 2023) learns to control the relative camera viewpoint using synthetic datasets, which helps generate new images with specified camera transformations, such as fixed camera viewpoints (such as left or right view) or randomly generated camera viewpoints. Despite being trained on synthetic datasets, Zero-1-to-3 still demonstrates strong zero-shot generalization capabilities to out-of-distribution datasets and real-world images. Figure 2 shows the different views generated by Zero-1-to-3 based on the input image.
4 Methodologies
4.1 Overview of ZeroVLM
To address the visual spatial reasoning task, we propose ZeroVLM, a model that leverages Zero-1-to-3 to infer various views of input images and employs a VLM to generate answers using the combined multiview images and our specially designed view prompts. The overview of our model is illustrated in Figure 3. In particular, ZeroVLM is a novel visual-language model that combines large language models (such as LLaMA (Touvron et al., 2023)) with high-level vision models, aiming to enhance visual-spatial reasoning capabilities by leveraging 3D reconstruction techniques. Given an image i, ZeroVLM uses Zero-1-to-3 to infer its various views $\hat{i}_{R,T}$ and performs image stitching from multiple views, thereby providing richer spatial information. The 3D reconstruction feature provided by Zero-1-to-3 can enable ZeroVLM to better understand and infer spatial relationships by viewing objects from different angles. At the same time, a specialized view prompt is designed into ZeroVLM to further enhance its visual spatial reasoning capabilities. This prompt helps guide the model to focus on relevant spatial views and improve its interpretation accuracy. The architecture of our ZeroVLM is shown in Figure 3.
4.2 Data Augmentation by 3D Reconstruction
In our work, we use Zero-1-to-3 for 3D reconstruction on the dataset to generate $\hat{i}_{R,T}$ from different viewpoints, e.g. the left, right and random views. Our investigation not only explores whether the creation of single-view images can enhance the visual spatial reasoning capabilities of VLMs but also examines whether multi-view images can help this improvement, where multi-view images are synthesized from different $\hat{i}_{R,T}$. Because multi-view images can provide richer spatial information, the model can get the opportunity to observe the same scene from different angles, thereby capturing a more comprehensive spatial layout and the relationship between objects, and different views can provide redundant information, so that the model can still make accurate judgments when facing noise or partial information loss. Therefore, in our work, after inferring different views of the input images, we further construct multi-view images from this single views by stitching them, and test their effectiveness against the single view images. Figure 2 shows an example of the multi-view image constructed by stitching multiple single view images generated through Zero-1-to-3.
By synthesizing these various multi-view images, we aim to determine their effectiveness in improving the spatial reasoning abilities of VLMs.
4.3 View Prompt
In our work, we first test whether 3D reconstruction of images can improve the accuracy of VLM’s visual spatial ability at the image level. To further explore whether the context prompt can enhance VLM’s visual spatial reasoning ability, we introduced a special prompt called view prompt in the experiment. This view prompt varies depending on the input image and its views. Figure 4 shows two view prompt examples over a single view and multiple views of the input.
We designed a variety of view prompts based on the content of different view images to guide VLM to better understand and reason about the view spatial relationship between the target object and other objects in the image. We first used Zero-1-to-3 (Liu et al., 2023c) to perform 3D reconstruction of VSR (Liu et al., 2023a) dataset and What’sUp (Kamath et al., 2023) dataset from different viewpoints. Then during the inference of VLM, we use a prompt consisting of the question, the view prompt, and the stitched view image, to generate the answer. Figure 5 provides a detailed description of the process.
Figure 4: These view prompts are manually constructed by us. View prompt comparison between single-view images and multi-view images. {question} is the corresponding question in the prompt.
Figure 5: For both single-view and multi-view datasets, we employed a view prompt. The aim was to explore whether perspective language models could enhance their visual spatial reasoning abilities through improvements at the textual level.
5 Experiments
5.1 Experimental Setup
Datasets. We selected two datasets for investigation: the Visual Spatial Reasoning (VSR) (Liu et al., 2023a) dataset and the What’sUp  (Kamath et al., 2023) dataset. The statistics of the two datasets are summarized in Table 1. The VSR dataset focuses on a wide range of spatial relations and linguistic phenomena, emphasizing various visual spatial relations and describing these relations. The What’sUp dataset focuses on household items and uses different prepositions to describe spatial relations. Both datasets cover a variety of visual spatial relations (such as “below”, “in front of”, “beside”, etc.).
DataSet Train Development Test Total
VSR (Random Split) 7680 1097 2195 10972
VSR (Zero shot) 4713 231 616 5560
What’sUp (SubSet-A) 200 110 111 421
What’sUp (SubSet-B) 200 108 100 408
Table 1: The number of images in various datasets.
Baselines. In our research, we employ LLaVA and MiniGPT-4 as the primary VLMs. LLaVA utilizes multi-modal language-image instruction data for instruction adjustment, employing CLIP-ViT-L-336px (Liu et al., 2023b) as the visual encoder and MLP projection as the visual language cross-modal connector, achieving comprehensive understanding of both visual and language inputs. The architecture of MiniGPT-4 involves aligning a frozen visual encoder with a frozen high-level language model, e.g. Vicuna (Chiang et al., 2023), through projection layers, ensuring correct alignment of visual features with the high-level large language model. MiniGPT-4 can exhibit advanced multi-modal capabilities similar to those of GPT-4 (Achiam et al., 2023).
Backbones. We use LLaVA and MiniGPT-4 as the backbone VLMs of our ZeroVLM in all the experiments. In particular, we denote ZeroVLM (L) as our model based on the LLaVA (Liu et al., 2024) model, which can process image inputs and improve efficiency and performance through joint learning of image and text instructions, while ZeroVLM (M) as our model based on the MiniGPT-4 (Zhu et al., 2023)) model, which combines the powerful generation ability of language models with visual information capabilities.
Task Setting. We classify visual spatial relationship questions for VLMs into two types using our dataset split, which is applied consistently across all four datasets. The first type disrupts the visual spatial relationship between the target object and another object (for example, the correct image description is “The apple is to the left of the banana, ” while our description is “The apple is above (below, to the right of, in front of, etc.) the banana”). The second type describes visual spatial relationships involving objects not present in the image but related to the target object (for example, the correct image description is “The apple is to the left of the banana, ” but our description is “The watermelon (an object that does not exist besides the apple) is above the banana”). These classifications are employed to evaluate the visual spatial reasoning abilities of VLMs and highlight the differences between our dataset and the original datasets.
Evaluation Metric. For evaluation, we judge the accuracy of the visual spatial reasoning ability of ZeroVLM (L) and ZeroVLM (M) based on the answers answered by ZeroVLM (L) and ZeroVLM (M). The accuracy of ZeroVLM (L) and ZeroVLM (M) responses determines their accuracy in identifying these relations. Since both datasets involve binary classification tasks, we use accuracy as the evaluation metric.
Implementation Details. We obtain new datasets through 3D reconstruction, consisting of single-view and multi-view datasets. In the single-view dataset, “Origin” refers to the original, unprocessed data, “Left” refers to the data processed by 3D reconstruction from the left viewpoint, and “Right” refers to the data processed by 3D reconstruction from the right viewpoint. “Random” refers to the data processed from a random viewpoint. In the multi-view dataset, there are two types of “Multi-view”, one is multi-view without original images, and the other is multi-view with original images. Multi-view without original images is composed of single-view images, while multi-view with original images is composed of original images plus different single-view images. Our goal is to explore whether the visual spatial reasoning ability of VLMs increases or decreases with the addition or omission of original images.
Model VSR (Random Split) VSR (Zero Shot) What’sUp (A) What’sUp (B)
Human 95.4 95.4 100 100
CLIP (frozen) 56.0 54.50 - -
CLIP (FT) 65.10 - - -
VisualBERT 55.20 51.00 - -
ViLT 69.30 63.00 - -
LXMERT 70.10 61.20 - -
ZeroVLM (M) 53.80 52.43 55.76 53.84
ZeroVLM (L) 70.29 70.94 71.74 80.76
Table 2: Accuracy performance of the compared models. Missing values indicate that ZeroVLM (L) and ZeroVLM (M) were not tested on the visual spatial reasoning on the VSR and What’sUp datasets. Bests are in bold, and second bests are underlined.
L-VR-VRa-VM-V404040404545454550505050555555556060606065656565707070707575757580808080858585859090909095959595100100100100Accuracy (%)(a) VSR (Random-Split)ZeroVLM (L)ZeroVLM (M)
L-VR-VRa-VM-V404040404545454550505050555555556060606065656565707070707575757580808080858585859090909095959595100100100100Accuracy (%)(b) VSR (Zero-Shot)ZeroVLM (L)ZeroVLM (M)
L-VR-VRa-VM-V404040404545454550505050555555556060606065656565707070707575757580808080858585859090909095959595100100100100Accuracy (%)(c) What’sUp (SubSet-A)ZeroVLM (L)ZeroVLM (M)
L-VR-VRa-VM-V404040404545454550505050555555556060606065656565707070707575757580808080858585859090909095959595100100100100Accuracy (%)(d) What’sUp (SubSet-B)ZeroVLM (L)ZeroVLM (M)
Figure 6: L-V represents the left-view images, R-V represents the right-view images, Ra-V represents the random-view images, and M-V represents the multi-view images (excluding the original image). Note that here we use ZeroVLM (L) and ZeroVLM (M) for fair comparison.
5.2 Experimental Result
Overall performance. Our experimental aims to comprehensively evaluate the visual spatial reasoning capabilities of VLMs ZeroVLM (L) and ZeroVLM (M) on various datasets. We seek to understand how different datasets (including original, single-view, and multi-view datasets) affect the performance of these models in visual-spatial reasoning tasks. We first test ZeroVLM (L) and ZeroVLM (M) on the original dataset (without 3D reconstruction) and compare the results with those of VSR (Liu et al., 2023a). The result of this comparison is reported in Table 2. Our experimental results show a baseline comparison of ZeroVLM (L) and ZeroVLM (M) testing the original dataset with VSR (Liu et al., 2023a), highlighting the basic performance of ZeroVLM (L) and ZeroVLM (M).
Additionally, we tested single-view 3D reconstructions from the VSR and What’sUp datasets, and multi-view datasets derived from these single views. We also assessed their visual spatial reasoning accuracy on single-view and multi-view datasets using different view prompts for ZeroVLM.
Single View vs Multiple View. Our study investigated the impact of single-view versus multi-view 3D reconstruction on the visual spatial reasoning capabilities of VLMs. We assessed two VLMs: ZeroVLM (L) and ZeroVLM (M), using both single-view and multi-view datasets. The results, shown in Figure 6, reveal that ZeroVLM excels in single-view conditions with an average accuracy of 79.28%, but its performance drops to 58.42% in multi-view conditions. ZeroVLM (M) performs less well overall, with average accuracy of 56.95% for single-view and 56.23% for multi-view. Overall, ZeroVLM (L) outperforms ZeroVLM (M) in visual spatial reasoning tasks, particularly in single-view scenarios. While ZeroVLM (M) possesses unique advantages in handling multi-modalities, there is still room for improvement in specific visual spatial reasoning tasks (Battaglia et al., 2018).
View Type View View Prompt VSR (Random Split) VSR (Zero Shot) What’sUp (A) What’sUp (B)
Origin $\times$ 70.29 70.94 71.14 80.76
L-V $\times$ 81.80 82.35 82.69 71.15
R-V $\times$ 78.60 79.41 80.76 78.84
Single-View Ra-V $\times$ 80.60 83.33 78.84 73.07
L-V $\checkmark$ 87.80 87.25 88.46 75.21
R-V $\checkmark$ 86.40 88.23 90.38 83.71
Ra-V $\checkmark$ 87.00 89.21 86.53 78.33
M-V $\times$ 55.60 60.78 59.61 57.69
Multi-View Origin + L-V $\checkmark$ 64.60 50.98 63.46 65.23
Origin + L-V + R-V $\checkmark$ 67.40 61.76 69.23 68.38
Origin + M-V $\checkmark$ 72.20 59.80 65.38 63.21
Table 3: Performance comparison of ZeroVLM (L) under different view combinations. L-V represents the left-view images, R-V represents the right-view images, Ra-V represents the random- view images, and M-V represents the multi-view images (excl. the original image). Bests are in bold, and second bests are underlined.
Effect on different view prompts. The primary goal of our concluding study was to explore the potential enhancement of visual spatial reasoning abilities in VLMs by employing different view prompts. We sought to investigate if leveraging contextual connections through textual cues could significantly improve the performance of VLMs in understanding and reasoning about visual spatial relationships. To achieve this, we applied various view prompts with ZeroVLM and assessed its visual spatial reasoning abilities across both single-view and multi-view datasets. These view prompts were designed to provide contextual cues that could aid the model in interpreting visual spatial information more effectively. The results of this comprehensive assessment are presented in Table 3. The experiment demonstrated that the use of different prompts can indeed enhance the visual spatial reasoning capabilities of VLMs. Overall, our study concludes that employing a view prompt can significantly enhance the visual spatial reasoning abilities of VLMs.
Case study. Our study aims to investigate the impact of 3D reconstruction on the visual spatial reasoning capabilities of VLM. We assess whether 3D reconstruction (Zero-1-to-3) enhances these capabilities from the original dataset to the single-view dataset and compare the performance of ZeroVLM (L) and ZeroVLM (M) using single-view images after 3D reconstruction. Comparative analysis focuses on the visual spatial reasoning capabilities of ZeroVLM on both datasets. Figure 7 illustrates the enhancement by ZeroVLM using the single-view dataset. Figure 8 presents the performance comparison between ZeroVLM (L) and ZeroVLM (M). Results show that ZeroVLM (L) has improved visual spatial reasoning with the single-view dataset after 3D reconstruction and outperforms ZeroVLM (M). This suggests that 3D reconstruction provides useful spatial information aiding visual reasoning.
Figure 7: Different results were obtained by ZeroVLM when evaluating datasets with and without 3D reconstruction. View prompts are omitted to save space.
Figure 8: Different results were obtained when inputting the 3D reconstructed datasets into ZeroVLM (L) and ZeroVLM (M). View prompts are omitted to save space.
6 Conclusion
This study presents a novel model, called ZeroVLM, for enhancing the visual spatial reasoning capability of VLMs through 3D reconstruction. In particular, ZeroVLM use the Zero-1-to-3 to generate different views over different angles of the input image, and generate answers with a specially designed view prompt. We validate the effectiveness of our VLMs by comparative experiments from various views, including single and multiple viewpoints, and evaluated the performance of LLaVA and MiniGPT-4 under our ZeroVLM. The experimental results suggest that 3D reconstruction with a view prompt from a single perspective can effectively optimize the model’s visual spatial reasoning ability. Future research could also develop models that dynamically adjust viewpoints based on task requirements and integrate additional modal information, such as video and audio, to enhance multi-modal processing capability.
7 Limitations
Although ZeroVLM demonstrates significant performance improvements, we still identify the following limitations of our work: (1) Data dependency: ZeroVLM used specific visual spatial reasoning datasets in our experiments. Although these datasets cover a variety of scenarios, they may not fully represent all the complex spatial relationships in the real world; (2) Diversity of datasets: Although we used multiple datasets for testing, these datasets may not fully cover all possible application scenarios. Therefore, the model may need further training and adjustment to ensure its generality and robustness when processing different types of images and tasks; (3) The generalization ability of the model: Although the Zero-1-to-3 model performs well in zero-shot generalization, in some extreme cases, the model may still not accurately capture the spatial relationship. Therefore, considering other methods for the 3D reconstruction module is one of our future work directions.
8 Potential risks
The development and deployment of VLMs for visual spatial reasoning present several potential risks. One significant risk is the dependency on specific datasets for training, which may not encompass the full diversity of spatial relationships encountered in real-world scenarios. This can lead to models that perform well in controlled environments but fail to generalize effectively. Additionally, the computational resources required for 3D reconstruction and multi-view generation are substantial, posing scalability and real-time application challenges. Another concern is the potential for bias in the datasets, which can result in models that unfairly represent or perform poorly on underrepresented spatial arrangements or object types. Furthermore, privacy and security issues arise when handling large amounts of visual data, necessitating strict adherence to data protection guidelines. Lastly, there are ethical considerations regarding the misuse of enhanced spatial reasoning capabilities for surveillance or other invasive applications, underscoring the need for transparency and accountability in the deployment of these technologies. Addressing these risks is essential to ensure the effective and responsible use of VLMs in visual spatial reasoning.
Appendix A Appendix
This section contains additional results. Figure 9 illustrates the differences between LLaVA and ZeroVLM (L). In our study, we prepare the required questions in JSON file format and modify ZeroVLM (L) to recognize and accept JSON file inputs. We enable ZeroVLM (L) to identify the corresponding questions by matching the input image names with the image names in the JSON file.Figure 10 illustrates the single-view and multi-view datasets obtained after 3D reconstruction of the VSR dataset and the What’sUp dataset.
Figure 9: The primary difference between LLaVA and ZeroVLM (L) lies in their input handling. LLaVA processes a single image as input, whereas ZeroVLM (L) is designed to handle multiple images. Due to the limited expressive capacity of individual images, we restrict the input for ZeroVLM (L) to two images.
Figure 10: The 3D reconstructed single-view and multi-view versions of each dataset are presented.